A=[3 0 2]

[4 -6 3]

[-2 1 8]


B=[-5 1 1]

[0 3 0]

[7 6 2]


C=[1 1 1]

[2 3 - 1]

[3 - 5 - 7]

Verify the following expressions(where possible and give reasons)

(i) A+(B+C) =(A+B) +C and A(BC) =(AB) C

(ii) (a-b) C=aC - bC and a(B - C) =aB - aC

(iii) (A^T) ^T=A and (A - B) ^T=A^T - B^T


1
Expert's answer
2021-06-01T09:13:55-0400

A = "\begin{bmatrix}
 3 & 0 & 2 \\
 4 & -6 & 3\\
 -2 & 1 & 8\\ 
\end{bmatrix}"


B = "\begin{bmatrix}
 -5 & 1 &1 \\
 0 & 3 & 0 \\
7 & 6 & 2 \\
\end{bmatrix}"


C = "\begin{bmatrix}
 1 & 1 &1\\
 2 &3&-1\\
3&-5&-7 
\end{bmatrix}"




(i) A+(B+C) =(A+B) +C and A(BC) =(AB) C


a) A+(B+C) =(A+B) +C


LHS:

= A+(B+C)


="\begin{bmatrix}
 3 & 0 & 2 \\
 4 & -6 & 3\\
 -2 & 1 & 8\\ 
\end{bmatrix}" + "\begin{bmatrix}
 -4 & 2 &2 \\
 2 & 6 & -1 \\
10 & 1 & -5 \\
\end{bmatrix}"


= "\begin{bmatrix}
 -1 & 2 &4\\
 6 & 0 &2\\
8&2&3
\end{bmatrix}"



RHS:

(A+B) +C


= "\begin{bmatrix}
-2&1&3\\
4&-3&3\\
5&7&10
\end{bmatrix}" + "\begin{bmatrix}
 1 & 1 &1\\
 2 &3&-1\\
3&-5&-7 
\end{bmatrix}"


= "\begin{bmatrix}
 -1 & 2 &4\\
 6 & 0 &2\\
8&2&3
\end{bmatrix}"



LHS = RHS




b) A(BC) =(AB) C

LHS:


= "\begin{bmatrix}
 3 & 0 & 2 \\
 4 & -6 & 3\\
 -2 & 1 & 8\\ 
\end{bmatrix}" * "\begin{bmatrix}
0&-7&-13\\
6&9&-3\\
25&15&-13\\








\end{bmatrix}"


= "\begin{bmatrix}
50&9&-65\\
39&-37&-73\\
206&143&-81\\
\end{bmatrix}"



RHS:


= "\begin{bmatrix}
 -1 & 15&7 \\
 1 & 4&10\\
66&49&14\\
\end{bmatrix}" * "\begin{bmatrix}
 1 & 1 &1\\
 2 &3&-1\\
3&-5&-7 
\end{bmatrix}"



= "\begin{bmatrix}
50&9&-65\\
39&-37&-73\\
206&143&-81\\
\end{bmatrix}"




LHS = RHS





(ii) (a-b) C=aC - bC and a(B - C) =aB - aC


a) (a-b) C=aC - bC

LHS:


= (a - b) * "\begin{bmatrix}
 1 & 1 &1\\
 2 &3&-1\\
3&-5&-7 
\end{bmatrix}"


= "\begin{bmatrix}
 a-b & a-b &a-b\\
 2(a - b) &3(a - b)&-1(a - b)\\
3(a - b)&-5(a - b)&-7(a - b)\\ 
\end{bmatrix}"



RHS:


= a * "\begin{bmatrix}
 1 & 1 &1\\
 2 &3&-1\\
3&-5&-7 
\end{bmatrix}" - b * "\begin{bmatrix}
 1 & 1 &1\\
 2 &3&-1\\
3&-5&-7 
\end{bmatrix}"


= "\begin{bmatrix}
 a & a &a\\
 2a &3a&-a\\
3a&-5a&-7a 
\end{bmatrix}" - "\begin{bmatrix}
 b & b &b\\
 2b &3b&-b\\
3b&-5b&-7b 
\end{bmatrix}"



= "\begin{bmatrix}
 a-b & a-b &a-b\\
 2(a - b) &3(a - b)&-1(a - b)\\
3(a - b)&-5(a - b)&-7(a - b)\\ 
\end{bmatrix}"



LHS = RHS


b) a(B - C) =aB - aC


LHS:


= a * "\begin{bmatrix}
-6&0&0\\
-2&0&1\\
4&11&9
\end{bmatrix}"


= "\begin{bmatrix}
-6a&0&0\\
-2a&0&a\\
4a&11a&9a
\end{bmatrix}"



RHS:


= aB - aC

= a * "\begin{bmatrix}
 -5 & 1 &1 \\
 0 & 3 & 0 \\
7 & 6 & 2 \\
\end{bmatrix}" - a * "\begin{bmatrix}
 1 & 1 &1\\
 2 &3&-1\\
3&-5&-7 
\end{bmatrix}"


= "\begin{bmatrix}
-6a&0&0\\
-2a&0&a\\
4a&11a&9a
\end{bmatrix}"


LHS = RHS






(iii) (A^T) ^T=A and (A - B) ^T=A^T - B^T


a) (A^T) ^T=A


LHS:


= ( AT ) T


= "(\begin{bmatrix}
 3 & 4 & -2 \\
 4 & -6 & 1\\
 -2 & 1 & 8\\ 
\end{bmatrix} )^T"


= "\begin{bmatrix}
 3 & 4 & -2 \\
 4 & -6 & 1\\
 -2 & 1 & 8\\ 
\end{bmatrix}" = A


LHS = RHS




(b) (A - B) ^T=A^T - B^T


LHS:


= "(\begin{bmatrix}
 8 & -1 & 1 \\
 4 & -9 & 3\\
 -9 & -5 & 6\\ 
\end{bmatrix} )^T"


= "\begin{bmatrix}
 8 & 4 & -9 \\
 -1 & -9 & -5\\
 1 & 3 & 6\\ 
\end{bmatrix}"




RHS:


= A^T - B^T

= "(\begin{bmatrix}
 3 & 0 & 2 \\
 4 & -6 & 3\\
 -2 & 1 & 8\\ 
\end{bmatrix} ) ^T" - "(\begin{bmatrix}
 -5 & 1 &1 \\
 0 & 3 & 0 \\
7 & 6 & 2 \\
\end{bmatrix})^T"


= "\begin{bmatrix}
 8 & 4 & -9 \\
 -1 & -9 & -5\\
 1 & 3 & 6\\ 
\end{bmatrix}"



LHS = RHS